OpenStudy (liizzyliizz):

Integral x^5 sec^2 (x^6) dx =

OpenStudy (anonymous):

Use substituion x^6=t. 6x^5dx=dt. x^5 dx= dt/6. So it is just integral of sec t dt /6.

OpenStudy (anonymous):

it is integral of sec ^2 t sorry.

OpenStudy (lgbasallote):

x^5dx is a perfect differential here...the answer is tan(x^6)/6

OpenStudy (lgbasallote):

oh wait...it's tan(x^6)/6 + C i forgot the arbitrary constant :D try checking it via derivatives.. it's correct :)